# out-of-domain OC20 dataset Open Catalyst 2020 dataset

### v1

@@
AQCat25: Unlocking spin-aware, high-fidelity machine learning potentials for heterogeneous catalysis Omar Allam, Brook Wander, Aayush R. Singh* SandboxAQ, Palo Alto, CA
@@ Introduction
Over the past three decades, computational approaches that couple first-principles density functional theory (DFT) with microkinetic modeling have become a cornerstone of modern heterogeneous catalysis research by providing a framework for rational catalyst design Nørskov et al. 2009; Greeley 2016; Motagamwala and Dumesic 2020; Chen et al. 2020; Xie et al. 2022. Numerous studies have linked atomic-scale surface chemistry to macroscopic kinetic observables, enabling the elucidation of complex reaction mechanisms for a variety of critical industrial heterogeneous catalytic processes including, but not limited to, ammonia synthesis Honkala et al. 2005; Singh et al. 2018, methanol synthesis Grabow and Mavrikakis 2011; Studt et al. 2012, Fischer-Tropsch synthesis Van Santen et al. 2013; Yao et al. 2019, selective hydrogenation Studt et al. 2008; He et al. 2021, steam reforming of methane Jones et al. 2008; Zhu and Yang 2021, the water-gas shift reaction Schumacher et al. 2005; Ghanekar et al. 2020, and ethylene epoxidation Stegelmann et al. 2004; Li et al. 2021. Many of these studies have leveraged unifying concepts such as d-band theory and the Brønsted-Evans-Polanyi relations Hammer and Nørskov 2000; Nørskov et al. 2002; Bligaard et al. 2004; Van Santen et al. 2009, which correlate the binding and transition-state energies of elementary reactions, facilitating the construction of volcano plots that predict optimal catalyst performance, in some cases even leading to experimentally validated discovery of new catalysts Besenbacher et al. 1998; Jacobsen et al. 2001. Despite these successes, the prohibitive cost of DFT largely limits its application to relatively simple networks of reactions taking place over idealized low-index facets of unary and binary catalyst materials Greeley 2016; Van Santen et al. 2009.
@@
Figure 1: A summary of the AQCat25 dataset and models. Spin polarization has been included for 12 important elements. A plane wave cutoff of 500 eV is used. Six new elements are added, when compared to the OC20 datasetChanussot et al. 2021 as well as 20 transition state adsorbates. Models were jointly tuned/trained from scratch on OC20 and AQCat25 to achieve good performance on their validation and test splits, respectively.
@@
Machine learning interatomic potentials (MLIPs) have emerged as an attractive alternative to estimate electronic structure properties at near-quantum accuracy for a small fraction of the computational cost Deringer et al. 2019; Unke et al. 2021; Pablo-García et al. 2023; Tang et al. 2024; Bozal-Ginesta et al. 2025; Yuan et al. 2025. These models learn the interactions required to predict the potential energy landscape of atomistic systems from large-scale public databases of DFT calculations Chen et al. 2021; Merchant et al. 2023; Batatia et al. 2023; Ko and Ong 2025. State-of-the-art models for heterogeneous catalysis are made possible by Meta FAIR’s Open Catalyst 20, 22, and 25 datasets Chanussot et al. 2021; Tran et al. 2023; Sahoo et al. 2025, which collectively consist of nearly 300 million single-point DFT calculations of adsorbate-surface interactions relevant for reactions of carbon, hydrogen, oxygen, and nitrogen over a diverse catalyst space spanning most of the periodic table. The introduction of machine learning methods into computational catalysis workflows has begun to enable studies of reaction network complexity Ulissi et al. 2017; Zhong et al. 2020; Margraf et al. 2023; Morandi et al. 2024 and catalyst structural dynamics Mou et al. 2023; Owen et al. 2024; Omranpour et al. 2025 that were completely inaccessible just 10 years ago.
@@
Although the sheer scale of these datasets has necessitated some compromises in the fidelity of the underlying training data, convergence of the resulting adsorption energies benefits from error cancellation and has been validated with respect to most DFT settings, with plane-wave cutoff and smearing width requiring some improvements to accurately capture total energies of non-metals Abdelmaqsoud et al. 2024. One of the most significant gaps in existing large-scale heterogeneous catalysis datasets is the treatment of magnetism. Since spin-polarized DFT calculations are considerably more expensive than spin-unpolarized calculations, spin is often omitted in the interest of scale and throughput Xu et al. 2025. The consequence of this choice is that the resulting models are not suitable for many industrially relevant catalytic processes such as ammonia synthesis Honkala et al. 2005 and Fischer-Tropsch synthesis Van Santen et al. 2013, which rely on on earth-abundant first-row transition metals (e.g., iron, cobalt, and nickel) that exhibit especially strong spin polarization effects on binding energies and activation barriers Sun et al. 2007a, b; Cao and Nørskov 2023; Zhang et al. 2024. As the field moves towards discovering new, low-cost, and sustainable catalytic materials to replace precious metals, the importance of treating magnetic effects when training foundational MLIPs becomes increasingly paramount.
@@
Alongside progress in data generation, developments in machine learning architectures, often based on equivariant graph neural networks, have improved performance on atomistic tasks. For heterogeneous catalysis, models like eSENFu et al. 2025, EquiformerV2 Liao et al. 2023 and EScAIP Qu and Krishnapriyan 2024 have achieved state-of-the-art results. A significant leap towards broader universality is the Universal Model for Atoms (UMA) Wood et al. 2025, trained on $\sim$500 million structures across diverse chemical domains (molecules, materials, catalysts). UMA modifies the eSEN architecture with additive embeddings for global context (charge, spin, DFT task) and a Mixture of Linear Experts (MoLE) routed by this context plus element composition. UMA’s core design goal is to accurately reproduce the original physics of each training task (e.g., spin-unpolarized OC20), operating as a multi-task surrogate rather than a model explicitly designed to perform cross-fidelity corrections between different levels of theory. A distinct advantage of total energy models like UMA is their ability to better capture, among other effects, restructuring of bare catalyst slabs Abdelmaqsoud et al. 2024.
@@
Other approaches have focused on integrating low- and high-fidelity data for related tasks, such as by augmenting node features with a fidelity one-hot encoding and applying both common and fidelity-specific weights in modified linear layers Kim et al. 2024 or utilizing a model’s intrinsic global state feature to embed fidelity context during message passing Ko and Ong 2025. For example, Ko and Ong demonstrated that a single multi-fidelity model trained with a small fraction of high-fidelity data could achieve similar accuracy to a single-fidelity model requiring eight times the amount of costly high-fidelity training data Ko and Ong 2025. Alternatively, other methods utilize architectural separation, such as dynamically using separate prediction heads branching from a shared backbone for each fidelity level Messerly et al. 2025; Gardner et al. 2025.
@@
While recent work has produced model architectures that can incorporate spin Xu et al. 2025; Chapman and Ma 2022; Deng et al. 2023; Anstine and Isayev 2023; Hu et al. 2023; Yang et al. 2024; Yu et al. 2024, their predictive power is limited by the absence of high-quality, spin-polarized training data for heterogeneous catalysis. Given the success of methods for adapting models to new data Tran et al. 2023; Musielewicz et al. 2022; Deng et al. 2025; Liu et al. 2025; King-Smith 2024; Radova et al. 2025, alongside advancements in training universal models from diverse datasets, we see a clear opportunity to develop improved foundational models specifically for spin-polarized, high-fidelity catalytic systems.
@@ Methods
DFT calculations were performed using the Vienna Ab Initio Simulation Package (VASP)Kresse and Hafner 1994; Kresse and Furthmüller 1996a, b; Kresse and Joubert 1999. A plane-wave cutoff energy of 500 eV was applied, and Gaussian smearing with a width of 0.1 eV was used. The revised Perdew-Burke-Ernzerhof (RPBE)Perdew et al. 1996; Zhang and Yang 1998 functional was chosen for its performance on heterogeneous catalyst systems, and the system’s geometry was optimized using the conjugate gradient algorithm. For systems containing Ce, Co, Cr, Cu, Fe, Mn, Mo, Ni, Os, Ru, V, or W, spin polarization was enabled to account for magnetic effects. For a full list of VASP parameters, please see the Supplementary Information.
@@
The AQCat25 bulk materials database was constructed by first updating and then expanding the OC20 dataset. Initially, the Materials Project (MP)Jain et al. 2013; Horton et al. 2025 database was queried for all structures containing only elements present in the OC20 dataset, subject to the constraints of a maximum of three unique elements per material and an energy above the convex hull ($E_{\text{hull}}$) of 0.1 eV/atom or less. Structures from the original OC20 dataset not found in this query were retained only if they were structurally unique. To expand the chemical space, a second query was performed using the same stability and size constraints but including six additional elements: Li, Ba, La, Ce, Mg, and F. The resulting set of new materials was then subsampled to ensure balanced representation. Up to 500 structures were randomly selected for each group containing a single new element, and up to 20 structures for each group containing a combination of new elements. The dataset was assembled by combining the updated OC20 dataset materials, the preserved unique structures, and the sampled new materials. Finally, the dataset was filtered to only contain bulks with up to 30 atoms per unit cell. Data splits were then assigned by attempting to preserve the original designations for all OC20 dataset materials and distributing new materials based on chemical composition to maintain consistency with the established OC20 dataset splitting methodology.
@@
The dataset is structured into three primary splits: in-domain (ID), out-of-domain (OOD) validation, and OOD test. Each OOD split is further categorized by the type of novelty introduced, either in the adsorbate or in the material slab. This strategy is designed to evaluate the model’s ability to generalize to novel systems it has not seen during training, in the same manner as the OC20 datasetChanussot et al. 2021. The ID split contains configurations where both the adsorbate and the material slab are present in the training set. The test and validation ID splits serve as a baseline for the model’s performance on familiar data and are sampled from the same distribution of the training set. Both OOD splits are designed to test the ability of machine learning models to generalize. The OOD validation set is used for hyperparameter tuning, while the OOD test set provides a final, unbiased evaluation of the model’s performance on unseen data.
@@
To sample high-energy states we took three approaches: (1) running molecular dynamics (MD) calculations, (2) placing transition state (TS) systems, and (3) rattling atoms. To sample high-energy states accessible at elevated temperatures, we performed MD calculations. Starting from a relaxed structure, we ran 80 steps of MD at 900 K. To provide the model with examples of highly distorted configurations relevant to chemical reactions, we extracted transition state structures from the OC20NEB datasetWander et al. 2025. These adsorbates were placed on new surfaces, followed by a short 5-step relaxation. This process generates data with high forces and energies, supporting the training of models that can handle reactive states. To further augment high-force data, we generated rattled configurations by randomly perturbing atomic positions. Two methods were used: (1) rattling all atoms and (2) rattling only adsorbate atoms, with displacements sampled from a normal distribution ($\sigma$ = 0.05, 0.1, 0.15, or 0.2 Å). Some rattled systems underwent a single DFT calculation, while others had a short 5-step relaxation. Systems whose max absolute force or absolute adsorption energy exceeds 50 eV/Å and 10 eV were excluded from training and evaluation.
@@
To explore the opportunity to train models with less costly DFT data, we considered data that include spin polarization but with settings that otherwise match the OC20 datasetChanussot et al. 2021. Notable differences between this data and the rest of the AQCat25 dataset are that we used a plane wave cutoff of 350 eV and Methfessel-Paxton smearing with a width of 0.2 eV. This data aids in understanding how the model handles the distinct physical regimes defined by fidelity and spin polarization. This dataset complements the existing high-fidelity spin-on/off (AQCat25) and spin-off OC20 data by filling a missing quadrant. Adsorption energies were computed using high-fidelity adsorbate references for all spin on systems. We found this to have little impact on the final target energies from preliminary tests.
We also wanted to form an understanding of model performance on the task of finding the minimum adsorption energy for an adsorbate-slab combination. To do this we constructed a small dense dataset, similar to the OC20dense dataset presented by Lan et al.Lan et al. 2023. For this dataset we selected 109 adsorbate-slab pairs. Adsorbates were selected to be disassociation reactants from the OC20NEB datasetWander et al. 2025. Slabs were selected randomly, but we selected the materials they were cut from more strategically. We included five unary materials, five binary non-metal materials, 46 binary intermetallics, 30 ternary intermetallics, and 23 ternary non-metals. Within these categories, the bulks were also randomly selected from the bulk database. For each adsorbate-slab pair, we performed 50 placements using the random site with heuristic placement mode in fairchem Shuaibi et al. 2025. These placements were relaxed with the same DFT settings as the broader AQCat25 dataset. The relaxed states were filtered using the same algorithms presented by Lan et al.Lan et al. 2023 to find desorption, dissociation, intercalation, and significant surface change.
@@
All systems were prepared using the publicly available fairchem package Shuaibi et al. 2025. Slab enumeration was performed using the underlying pymatgenSun and Ceder 2013; Tran et al. 2016 algorithm. Adsorbate placement was performed heuristically at random sites. Rattled systems were perturbed after adsorbate placement using the rattle functionality in ASEBahn and Jacobsen 2002; Larsen et al. 2017. For TS systems, they were placed as normal adsorbed intermediates would be by preparing a new adsorbate database with TS entries.
@@
A challenge in this work is training a single MLIP that can accurately predict energies and forces across a dataset containing multiple DFT settings. The combined training data spans four distinct physical regimes: high-fidelity spin-on, high-fidelity spin-off, low-fidelity spin-on, and the original low-fidelity spin-off. We therefore explored methods to introduce this DFT context, namely spin treatment and calculation fidelity, directly to the model. Inspired by the effectiveness of Feature-wise Linear Modulation (FiLM) Perez et al. 2018 and similar successful techniques Wood et al. 2025 for multi-task learning, the baseline models presented in this paper focus on adapting the EquiformerV2 (EV2) architectureLiao et al. 2023 using this approach. We did not focus on fine-tuning UMAWood et al. 2025 models because of licensing issues, but we expect those to have even better performance than the models presented here. The Feature-wise Linear Modulation (FiLM) technique Perez et al. 2018 provides an expressive conditioning mechanism. Rather than adding context via feature concatenation, FiLM applies a learned, feature-wise affine transformation ($\gamma F+\beta$) that can scale, shift, or suppress activations. This strategy of deep, additive modulation is similar in principle to the additive embedding mechanism successfully employed by the UMA family of models Wood et al. 2025. To evaluate performance on AQCat25 while retaining knowledge from OC20, we used the EV2Liao et al. 2023 model architecture with three variants and three training protocols. The variants were: (i) EV2 (unmodified), (ii) EV2-inFiLM, which applies additive FiLM Perez et al. 2018 shifts to the scalar (l=0l{=}0) channels at the input, and (iii) EV2-in+midFiLM, which applies the same modulation at the input and after each equivariant block. The protocols were: direct fine-tuning of OC20-pretrained checkpoints, cotuning those checkpoints with OC20 replay, and cotraining from scratch on mixed data from both AQCat25 and OC20.
@@
EV2 is an E(3)-equivariant transformer over atomic graphs: atoms form nodes, edges use pairwise distances and spherical harmonics, attention layers are equivariant, and feed-forward layers use $S^{2}$ activations Cohen et al. 2018; Zitnick et al. 2022; Passaro and Zitnick 2023; Liao et al. 2023. Energies are predicted by a scalar head, and forces are predicted directly via a vector head. Architectural hyperparameters are listed in Table 7.
@@
Figure 3: Summary statistics on the AQCat25 dataset and some comparisons to the OC20 datasetChanussot et al. 2021. (a) Element counts showing the frequency with which each element appears in the dataset. (b) The proportion of systems that fit into four material-type categories for the entire training splits OC20 dataset and AQCat25. (c & d) The distribution of adsorption energies and maximum forces across the training split of AQCat25 and the 2M training subsplit of OC20.
Some summary statistics about the AQCat25 dataset and how it compares to the OC20 dataset are shown in Figure 3. As can be seen in Figure 3b, there is a significantly higher proportion of non-metal systems and lower proportion of metal systems in the AQCat25 dataset compared to the OC20 dataset. The proportion of the other two categories, however, (non-metal & metalloid and metalloid) are roughly equal. Because non-metal systems typically have poorer performance compared to intermetallicsKolluru et al. 2022, we will look at key model performance metrics split over these material categories. The adsorption energy and maximum force distributions (Figure 3c-d) reveal that the AQCat25 dataset is biased towards higher force, higher energy systems when compared to OC20. This can be explained by the more aggressive approach taken when sampling high-force systems. Here, we used larger standard deviations to sample rattled configurations and also included the high energy transition state like systems.
@@
We wanted to explore opportunities to improve our data generation strategy to maximize model performance while minimizing the cost associated with dataset generation and model fine-tuning. To do this, we looked at the change in model performance as a function of two variables: number of DFT single points seen per system and number of slabs. For heterogeneous catalyst systems, there are two types of diversity the models must generalize across: (1) the adsorbates and (2) the material surfaces the adsorbates are adsorbed to. The latter is much more complex. We initially adopted a scheme of performing four adsorbate-slab relaxations per slab to reduce the number of slab relaxations that needed to be performed as this data is not directly used in model training for referenced energy models. This turned out to be a suboptimal choice, however, because material diversity is important to tackle. For the future, we would choose to perform one adsorbate-slab calculation per slab. In Figure 4, we show the relationship between model force performance, number of slabs seen, and number of DFT single points seen per system. Models were directly tuned starting from the publicly available 31M parameter EV2 model (All + MD) Liao et al. 2023.

### v2

@@
AQCat25: Unlocking spin-aware, high-fidelity machine learning potentials for heterogeneous catalysis Omar Allam1,†, Brook Wander1,†, SungYeon Kim1, Rudi Plesch2, Tyler Sours2, Jia-Min Chu2, Thomas Ludwig2, Jiyoon Kim2, Rodrigo Wang2, Shivang Agarwal2, Alan Rask2, Alexandre Fleury2, Chuhong Wang2, Andrew Wildman2, Thomas Mustard2, Kevin Ryczko2, Paul Abruzzo3, AJ Nish3, Aayush R. Singh1,* 1SandboxAQ Core Team, Palo Alto, CA
2SandboxAQ Support Team, Palo Alto, CA
3Nvidia Corporation, Santa Clara, CA
@@ Introduction
Over the past three decades, computational approaches that couple first-principles density functional theory (DFT) with microkinetic modeling have become a cornerstone of modern heterogeneous catalysis research by providing a framework for rational catalyst design norskov2009towards; greeley2016theoretical; motagamwala2020microkinetic; chen2020computational; xie2022achieving. Numerous studies have linked atomic-scale surface chemistry to macroscopic kinetic observables, enabling the elucidation of complex reaction mechanisms for a variety of critical industrial heterogeneous catalytic processes including, but not limited to, ammonia synthesis honkala2005ammonia; singh2018computational, methanol synthesis grabow2011mechanism; studt2012co, Fischer-Tropsch synthesis van2013mechanism; yao2019quantitative, selective hydrogenation studt2008identification; he2021catalyst, steam reforming of methane jones2008first; zhu2021first, the water-gas shift reaction schumacher2005trends; ghanekar2020catalysis, and ethylene epoxidation stegelmann2004microkinetic; li2021understanding. Many of these studies have leveraged unifying concepts such as d-band theory and the Brønsted-Evans-Polanyi relations hammer2000theoretical; norskov2002universality; bligaard2004bronsted; van2009reactivity, which correlate the binding and transition-state energies of elementary reactions, facilitating the construction of volcano plots that predict optimal catalyst performance, in some cases even leading to experimentally validated discovery of new catalysts besenbacher1998design; jacobsen2001catalyst. Despite these successes, the prohibitive cost of DFT largely limits its application to relatively simple networks of reactions taking place over idealized low-index facets of unary and binary catalyst materials greeley2016theoretical; van2009reactivity.
@@
Figure 1: A summary of the AQCat25 dataset and models. Spin polarization has been included for 12 important elements. A plane wave cutoff of 500 eV is used. Six new elements are added, when compared to the OC20 datasetchanussot2021open as well as 20 transition state adsorbates. Models were jointly tuned/trained from scratch on OC20 and AQCat25 to achieve good performance on their validation and test splits, respectively.
@@
Machine learning interatomic potentials (MLIPs) have emerged as an attractive alternative to estimate electronic structure properties at near-quantum accuracy for a small fraction of the computational cost deringer2019machine; unke2021machine; pablo2023fast; tang2024machine; bozal2025developing; yuan2025foundation. These models learn the interactions required to predict the potential energy landscape of atomistic systems from large-scale public databases of DFT calculations chen2021learning; merchant2023scaling; batatia2023foundation; ko2025data. State-of-the-art models for heterogeneous catalysis are made possible by Meta FAIR’s Open Catalyst 20, 22, and 25 datasets chanussot2021open; tran2023open; sahoo2025open, which collectively consist of nearly 300 million single-point DFT calculations of adsorbate-surface interactions relevant for reactions of carbon, hydrogen, oxygen, and nitrogen over a diverse catalyst space spanning most of the periodic table. The introduction of machine learning methods into computational catalysis workflows has begun to enable studies of reaction network complexity ulissi2017address; zhong2020accelerated; margraf2023exploring; morandi2024foundational and catalyst structural dynamics mou2023bridging; owen2024atomistic; omranpour2025machine that were completely inaccessible just 10 years ago.
@@
Although the sheer scale of these datasets has necessitated some compromises in the fidelity of the underlying training data, convergence of the resulting adsorption energies benefits from error cancellation and has been validated with respect to most DFT settings, with plane-wave cutoff and smearing width requiring some improvements to accurately capture total energies of non-metals abdelmaqsoud2024investigating. One of the most significant gaps in existing large-scale heterogeneous catalysis datasets is the treatment of magnetism. Since spin-polarized DFT calculations are considerably more expensive than spin-unpolarized calculations, spin is often omitted in the interest of scale and throughput xu2025spin. The consequence of this choice is that the resulting models are not suitable for many industrially relevant catalytic processes such as ammonia synthesis honkala2005ammonia and Fischer-Tropsch synthesis van2013mechanism, which rely on earth-abundant first-row transition metals (e.g., iron, cobalt, and nickel) that exhibit especially strong spin polarization effects on binding energies and activation barriers sun2007spin; sun2007spinbenz; cao2023spin; zhang2024spin. As the field moves towards discovering new, low-cost, and sustainable catalytic materials to replace precious metals, the importance of treating magnetic effects when training foundational MLIPs becomes increasingly paramount.
@@
Alongside progress in data generation, developments in machine learning architectures, often based on equivariant graph neural networks, have improved performance on atomistic tasks. For heterogeneous catalysis, models like eSENfu2025learning, EquiformerV2 liao2023equiformerv2 and EScAIP qu2024importance have achieved state-of-the-art results. A significant leap towards broader universality is the Universal Model for Atoms (UMA) wood2025family, trained on $\sim$500 million structures across diverse chemical domains (molecules, materials, catalysts). UMA modifies the eSEN architecture with additive embeddings for global context (charge, spin, DFT task) and a Mixture of Linear Experts (MoLE) routed by this context plus element composition. UMA’s core design goal is to accurately reproduce the original physics of each training task (e.g., spin-unpolarized OC20), operating as a multi-task surrogate rather than a model explicitly designed to perform cross-fidelity corrections between different levels of theory. A distinct advantage of total energy models like UMA is their ability to better capture, among other effects, restructuring of bare catalyst slabs abdelmaqsoud2024investigating.
@@
Other approaches have focused on integrating low- and high-fidelity data for related tasks, such as by augmenting node features with a fidelity one-hot encoding and applying both common and fidelity-specific weights in modified linear layers kim2024dataefficient or utilizing a model’s intrinsic global state feature to embed fidelity context during message passing ko2025data. For example, Ko and Ong demonstrated that a single multi-fidelity model trained with a small fraction of high-fidelity data could achieve similar accuracy to a single-fidelity model requiring eight times the amount of costly high-fidelity training data ko2025data. Alternatively, other methods utilize architectural separation, such as dynamically using separate prediction heads branching from a shared backbone for each fidelity level messerly2025multifidelity; gardner2025understanding.
@@
While recent work has produced model architectures that can incorporate spin xu2025spin; chapman2022machine; deng2023chgnet; anstine2023machine; hu2023spin; yang2024deep; yu2024spin, their predictive power is limited by the absence of high-quality, spin-polarized training data for heterogeneous catalysis. Given the success of methods for adapting models to new data tran2023open; musielewicz2022finetuna; deng2025systematic; liu2025fine; king2024transfer; radova2025fine, alongside advancements in training universal models from diverse datasets, we see a clear opportunity to develop improved foundational models specifically for spin-polarized, high-fidelity catalytic systems.
@@ Methods
DFT calculations were performed using the Vienna Ab Initio Simulation Package (VASP)kresse1994ab; kresse1996efficiency; kresse1996efficient; kresse1999ultrasoft. A plane-wave cutoff energy of 500 eV was applied, and Gaussian smearing with a width of 0.1 eV was used. The revised Perdew-Burke-Ernzerhof (RPBE)perdew1996generalized; zhang1998comment functional was chosen for its performance on heterogeneous catalyst systems, and the system’s geometry was optimized using the conjugate gradient algorithm. For systems containing Ce, Co, Cr, Cu, Fe, Mn, Mo, Ni, Os, Ru, V, or W, spin polarization was enabled to account for magnetic effects. For a full list of VASP parameters, please see the Supplementary Information.
@@
The AQCat25 bulk materials database was constructed by first updating and then expanding the OC20 dataset. Initially, the Materials Project (MP)jain2013commentary; horton2025accelerated database was queried for all structures containing only elements present in the OC20 dataset, subject to the constraints of a maximum of three unique elements per material and an energy above the convex hull ($E_{\text{hull}}$) of 0.1 eV/atom or less. Structures from the original OC20 dataset not found in this query were retained only if they were structurally unique. To expand the chemical space, a second query was performed using the same stability and size constraints but including six additional elements: Li, Ba, La, Ce, Mg, and F. The resulting set of new materials was then subsampled to ensure balanced representation. Up to 500 structures were randomly selected for each group containing a single new element, and up to 20 structures for each group containing a combination of new elements. The dataset was assembled by combining the updated OC20 dataset materials, the preserved unique structures, and the sampled new materials. Finally, the dataset was filtered to only contain bulks with up to 30 atoms per unit cell. Data splits were then assigned by attempting to preserve the original designations for all OC20 dataset materials and distributing new materials based on chemical composition to maintain consistency with the established OC20 dataset splitting methodology.
@@
The dataset is structured into three primary splits: in-domain (ID), out-of-domain (OOD) validation, and OOD test. Each OOD split is further categorized by the type of novelty introduced, either in the adsorbate or in the material slab. This strategy is designed to evaluate the model’s ability to generalize to novel systems it has not seen during training, in the same manner as the OC20 datasetchanussot2021open. The ID split contains configurations where both the adsorbate and the material slab are present in the training set. The test and validation ID splits serve as a baseline for the model’s performance on familiar data and are sampled from the same distribution of the training set. Both OOD splits are designed to test the ability of machine learning models to generalize. The OOD validation set is used for hyperparameter tuning, while the OOD test set provides a final, unbiased evaluation of the model’s performance on unseen data.
@@
To sample high-energy states we took three approaches: (1) running molecular dynamics (MD) calculations, (2) placing transition state (TS) systems, and (3) rattling atoms. To sample high-energy states accessible at elevated temperatures, we performed MD calculations. Starting from a relaxed structure, we ran 80 steps of MD at 900 K. To provide the model with examples of highly distorted configurations relevant to chemical reactions, we extracted transition state structures from the OC20NEB datasetwander2025cattsunami. These adsorbates were placed on new surfaces, followed by a short 5-step relaxation. This process generates data with high forces and energies, supporting the training of models that can handle reactive states. To further augment high-force data, we generated rattled configurations by randomly perturbing atomic positions. Two methods were used: (1) rattling all atoms and (2) rattling only adsorbate atoms, with displacements sampled from a normal distribution ($\sigma$ = 0.05, 0.1, 0.15, or 0.2 Å). Some rattled systems underwent a single DFT calculation, while others had a short 5-step relaxation. Systems whose max absolute force or absolute adsorption energy exceeds 50 eV/Å and 10 eV were excluded from training and evaluation.
@@
To explore the opportunity to train models with less costly DFT data, we considered data that include spin polarization but with settings that otherwise match the OC20 datasetchanussot2021open. Notable differences between this data and the rest of the AQCat25 dataset are that we used a plane wave cutoff of 350 eV and Methfessel-Paxton smearing with a width of 0.2 eV. This data aids in understanding how the model handles the distinct physical regimes defined by fidelity and spin polarization. This dataset complements the existing high-fidelity spin-on/off (AQCat25) and spin-off OC20 data by filling a missing quadrant. Adsorption energies were computed using high-fidelity adsorbate references for all spin on systems. We found this to have little impact on the final target energies from preliminary tests.
We also wanted to form an understanding of model performance on the task of finding the minimum adsorption energy for an adsorbate-slab combination. To do this we constructed a small dense dataset, similar to the OC20dense dataset presented by Lan et al.lan2023adsorbml. For this dataset we selected 109 adsorbate-slab pairs. Adsorbates were selected to be disassociation reactants from the OC20NEB datasetwander2025cattsunami. Slabs were selected randomly, but we selected the materials they were cut from more strategically. We included five unary materials, five binary non-metal materials, 46 binary intermetallics, 30 ternary intermetallics, and 23 ternary non-metals. Within these categories, the bulks were also randomly selected from the bulk database. For each adsorbate-slab pair, we performed 50 placements using the random site with heuristic placement mode in fairchem fairchem. These placements were relaxed with the same DFT settings as the broader AQCat25 dataset. The relaxed states were filtered using the same algorithms presented by Lan et al.lan2023adsorbml to find desorption, dissociation, intercalation, and significant surface change.
@@
All systems were prepared using the publicly available fairchem package fairchem. Slab enumeration was performed using the underlying pymatgensun2013efficient; tran2016surface algorithm. Adsorbate placement was performed heuristically at random sites. Rattled systems were perturbed after adsorbate placement using the rattle functionality in ASEase_old; ase-paper. For TS systems, they were placed as normal adsorbed intermediates would be by preparing a new adsorbate database with TS entries.
@@
A challenge in this work is training a single MLIP that can accurately predict energies and forces across a dataset containing multiple DFT settings. The combined training data spans four distinct physical regimes: high-fidelity spin-on, high-fidelity spin-off, low-fidelity spin-on, and the original low-fidelity spin-off. We therefore explored methods to introduce this DFT context, namely spin treatment and calculation fidelity, directly to the model. Inspired by the effectiveness of Feature-wise Linear Modulation (FiLM) perez2018film and similar successful techniques wood2025family for multi-task learning, the baseline models presented in this paper focus on adapting the EquiformerV2 (EV2) architectureliao2023equiformerv2 using this approach. We did not focus on fine-tuning UMAwood2025family models because of licensing issues, but we expect those to have even better performance than the models presented here. The Feature-wise Linear Modulation (FiLM) technique perez2018film provides an expressive conditioning mechanism. Rather than adding context via feature concatenation, FiLM applies a learned, feature-wise affine transformation ($\gamma F+\beta$) that can scale, shift, or suppress activations. This strategy of deep, additive modulation is similar in principle to the additive embedding mechanism successfully employed by the UMA family of models wood2025family. To evaluate performance on AQCat25 while retaining knowledge from OC20, we used the EV2liao2023equiformerv2 model architecture with three variants and three training protocols. The variants were: (i) EV2 (unmodified), (ii) EV2-inFiLM, which applies additive FiLM perez2018film shifts to the scalar (l=0l{=}0) channels at the input, and (iii) EV2-in+midFiLM, which applies the same modulation at the input and after each equivariant block. The protocols were: direct fine-tuning of OC20-pretrained checkpoints, cotuning those checkpoints with OC20 replay, and cotraining from scratch on mixed data from both AQCat25 and OC20.
@@
EV2 is an E(3)-equivariant transformer over atomic graphs: atoms form nodes, edges use pairwise distances and spherical harmonics, attention layers are equivariant, and feed-forward layers use $S^{2}$ activations cohen2018spherical; zitnick2022spherical; passaro2023reducing; liao2023equiformerv2. Energies are predicted by a scalar head, and forces are predicted directly via a vector head. Architectural hyperparameters are listed in Table 7.
@@
Figure 3: Summary statistics on the AQCat25 dataset and some comparisons to the OC20 datasetchanussot2021open. (a) Element counts showing the frequency with which each element appears in the dataset. (b) The proportion of systems that fit into four material-type categories for the entire training splits OC20 dataset and AQCat25. (c & d) The distribution of adsorption energies and maximum forces across the training split of AQCat25 and the 2M training subsplit of OC20.
Some summary statistics about the AQCat25 dataset and how it compares to the OC20 dataset are shown in Figure 3. As can be seen in Figure 3b, there is a significantly higher proportion of non-metal systems and lower proportion of metal systems in the AQCat25 dataset compared to the OC20 dataset. The proportion of the other two categories, however, (non-metal & metalloid and metalloid) are roughly equal. Because non-metal systems typically have poorer performance compared to intermetallicskolluru2022open, we will look at key model performance metrics split over these material categories. The adsorption energy and maximum force distributions (Figure 3c-d) reveal that the AQCat25 dataset is biased towards higher force, higher energy systems when compared to OC20. This can be explained by the more aggressive approach taken when sampling high-force systems. Here, we used larger standard deviations to sample rattled configurations and also included the high energy transition state like systems.
@@
We wanted to explore opportunities to improve our data generation strategy to maximize model performance while minimizing the cost associated with dataset generation and model fine-tuning. To do this, we looked at the change in model performance as a function of two variables: number of DFT single points seen per system and number of slabs. For heterogeneous catalyst systems, there are two types of diversity the models must generalize across: (1) the adsorbates and (2) the material surfaces the adsorbates are adsorbed to. The latter is much more complex. We initially adopted a scheme of performing four adsorbate-slab relaxations per slab to reduce the number of slab relaxations that needed to be performed as this data is not directly used in model training for referenced energy models. This turned out to be a suboptimal choice, however, because material diversity is important to tackle. For the future, we would choose to perform one adsorbate-slab calculation per slab. In Figure 4, we show the relationship between model force performance, number of slabs seen, and number of DFT single points seen per system. Models were directly tuned starting from the publicly available 31M parameter EV2 model (All + MD) liao2023equiformerv2.